OpenStudy (anonymous):

what is the improper fraction of 12/5

jimthompson5910 (jim_thompson5910):

Do you want to convert this to a mixed number?

OpenStudy (anonymous):

if you do its 2 2/5

OpenStudy (solomonzelman):

12/5 is an improper fraction. When the numerator is greater than the denominator \(\large\color{black}{ \rm mixed~~number~~~~ 5 \frac{2}{3}= \frac{(5 \times 3)+2}{3}=\frac{17}{3}~~~~improper~~fraction. }\)

OpenStudy (solomonzelman):

I showed an example with 5 and 2/3